Suppose your broker give you the following information:

Spot exchange rate (USD/EUR) = 1.1370

One year forward rate (USD/EUR) = 1.1405

One year USD interest rate = 0.87%

One year Euro interest rate = 0.65%

a. Is there any violation of interest rate parity?

b. How would you take advantage of any arbitrage situation?

c. What is your profit?

d. Suggest an equilibrium value for the forward rate
Answer #1

a. Yes there is violation of interest rate parity

b. Arbitrage Situation

Borrow 1 EUR for 1 year @ 0.65%

Present Value of borrowing = 1/1.0065 = 0.993541

Convert 0.993541 EUR in spot market in USD , Spot rate USD/EUR = 1.1370

0.993541 EUR = 0.993541/1.1370 = 0.873827USD

Invest 0.873827USD @ 0.87% interest rate for 1 year

Value after 1 year = 0.873827(1.0087) = 0.881429 USD

Convert 0.881429 USD into EUR , one year forward rate USD/EUR = 1.1405

0.881429 USD = 0.881429*1.1405 = 1.00527 EUR

Payback 1 EUR

c. Profit = 1.00527EUR - 1EUR = 0.00527

d. Equilibrium value of forward rate USD/EUR= 1/0.881429 = 1.13452